# for Academic Information Seeking

### v1

@@ 9 Conclusion
In this work, we propose an API-Using methodology, SoAy, to apply LLMs to academic information searching. The core technique of SoAy is the high-quality automatic built dataset, SoAyBench, providing 3960 academic triplets data. We align LLMs to generate codes on pre-constructed guidance, which could face the coupling and execution logic diverse challenges. An evaluation strategy, SoAyEval, is proposed to evaluate the AMiner API-using capability. Experiments show SoAy’s effectiveness, efficiency, and usability. In the future, we will continue to enlarge the SoAyBench and apply SoAy to more scenarios and domains.
10 Limitations and Future Directions
In this paper, we propose an approach that combines automated data generation with LLM alignment for academic query API-based tasks and demonstrates its effectiveness. The released SoAyBench and the proposed SoAyEval could support for evaluating related methods in subsequent studies. However, there are still some worthwhile explorations that we have not conducted:
  • SoAyBench is a large-scale dataset constructed through automation, consisting of 3960 triplet data instances. Although we incorporate quality checks in the solution library construction, query generation, and code generation processes, and have human experts inspect and refine the test set, we must acknowledge that we cannot guarantee the provided tests in SoAyBench align perfectly with people’s expectations for the API system. Some prior worksQin et al. (2023b) validate the quality of their benchmarks by comparing the results of tested methods on their own benchmark with manually evaluated results. This is an aspect we plan to further explore and validate in the future, which may lead to more human-aligned automated constructed datasets.
  • The API-using approach we propose addresses the task of guiding large models to utilize complex academic search APIs. However, we have not yet validated whether our method has advantages in simpler scenarios. We will continue to expand SoAyBench and conduct practicality and effectiveness validations in various scenarios.
  • In practical applications, we have encountered an inconvenient point of updating the deployed system: if we want to incorporate new APIs on top of existing ones, especially when adding a large number of APIs, the context-guided SoAyGPT needs to adjust the structure of the prompt within a limited input space. This poses a challenge that requires a trade-off between the breadth and granularity of examples. Although our proposed training-based SoAyLLaMA theoretically allows for the unlimited expansion of APIs by adding sufficient training data, this is an aspect we have not explored yet. In the future, we plan to explore methods for continuous learning with SoAyLLaMA to enable rapid inclusion of new API data without compromising the model’s mastery of existing APIs.
In conclusion, the quality of data generation, scalability, and the scenarios presented in this paper opens up numerous possibilities for further exploration. In the future, we will continue to fulfill the current work and explore further in these mentioned directions.
11 Acknowledgements
@@
This project would like to thank the data annotators at Zhipu AI for their help and support. All GPU compute and API costs are provided by Zhipu AI.
12 Ethical Considerations
12.1 New datasets
Using SoAy, our new dataset is built on the AMiner dataset. The construction of this new dataset relies on the efforts of annotators. We value their contributions and provide them with reasonable compensation based on the quantity of annotations and the duration of their work. For the characteristics of the dataset, it focuses on the academic field and is suitable for academic researchers. Furthermore, as talked before, to prevent additional risks, we establish an evaluation benchmark, which is further enhanced by additional human verification to ensure the quality of this new dataset and its application effectiveness.
12.2 NLP application
The intended use of the SoAy trained LLMs is to facilitate researchers in better accessing academic literature and other related resources. It allows for querying in natural language, reducing the technical barriers for scholars and others with academic information retrieval needs. However, just like the issues inherent to LLMs, sometimes it may provide incorrect information, potentially leading to misguidance for users, which may be a serious disruption for researchers since academic information needs to be accurate.
@@ 12.2 NLP application
In the appendix, we’ll present in three folds the details of SoAy and SoAyBench, along with more experimental details and conclusions. Part I focus on the implementation of SoAy, including the details of SoAyGPT in  A, SoAyLLaMA in  B and SoAyBench in  C. Part II shows more about the experiments, including the online evaluation in  4, the implementation of SoAy on other academic information seeking systems in  F, and more interesting findings from the results on SoAyBench in  E. Part III lists all the prompts used in SoAy and SoAyGPT.
Part I. Implementation Details
The content of this section primarily introduces some implementation details of the SoAy method, including the inference architecture of SoAyGPT, the training details of SoAyLLaMA, as well as the data details and API environment setup details of SoAyBench. We implement SoAyGPT on three distinct versions of GPT, specifically 3.5: gpt-3.5-turbo-0613; 3.5-16k: gpt-3.5-turbo-16k-0613, and 4: gpt-4-0613. We conduct the fine-tuning on 3 different variants of the LLaMA model, respectively Chat-7B: Llama-2-7b-chat-hf, Code-7B: CodeLlama-7b-Instruct-hf, and Code-13B: CodeLlama-13b-Instruct-hf.
Appendix A SoAyGPT
@@
SoAyGPT is an alignment method designed for a series of large-scale models that are closed-source yet possess high command-following capabilities. Specifically, SoAyGPT employs in-context learning to construct three LLM agents Wang et al. (2024) (Solution Retrieval Agent, Code Generation Agent, and Answer Generation Agent) and an environment operable by these agents to complete the alignment process. The agents and environments are designed:
Solution Retrieval Agent. Given the user input $Q$, the solution retrieval agent obtains the appropriate solution from the solution library $S$. It’s important to note that the retrieval process here diverges from traditional information retrieval functions, which extract the exact required element from a set. Essentially, this retrieval module presents all possible solutions as the prompt to the LLMs, enabling them to generate a suitable solution. Consequently, it’s possible that the produced solution differs from any solution in $L_{S}$.
Code Generation Agent. The code generation agent must generate the correct program under the guidance of the retrieved solution $S$. Since the solution might not be included in $L_{S}$, if the retrieved solution $S$ is in $L_{S}$, we simply use $P_{p}$ specifically defined for $S$ to generate the required program. Otherwise, SoAy has to use $P_{g}$ that is defined for all the possible solutions to guide the LLMs in generating a program, as no specific solution could be identified. SoAy employs Python as the programming language for each program, allowing it to incorporate not only sequential but also branching and looping logic.
Code Execution Environment. SoAy employs a virtual Python environment as the executor to obtain the results, serving as the knowledge $K$. Here, $K$ is expected to be the exact answer to $Q$ if $Q$ asks about some precise information, like the number of citations for a scholar. Otherwise, $K$ provides helpful information if $Q$ has an ambiguous intent.
Answer Generation Agent. Finally, SoAy generates a natural language answer $A$ based on $Q$ and $K$, guided by $P_{a}$.
Appendix B SoAyLLaMA
@@
SoAyLLaMA is designed for supervised fine-tuning of open-source models when training resources are available. Through the training stage, SoAyLLaMA integrates an API description and queries as inputs and combines solutions and programs as responses. During the inference stage, SoAyLLaMA concurrently completes the Planning and Formatting processes to derive solutions and code. Similar to SoAyGPT, it utilizes the environment to execute and obtain answers based on $K$. We conduct the fine-tuning on 3 different variants of LLaMA Model, respectively Chat-7B: Llama-2-7b-chat-hf [^2] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf, Code-7B: CodeLlama-7b-Instruct-hf [^3] https://www.modelscope.cn/models/AI-ModelScope/CodeLlama-7b-hf/summary, Code-13B: CodeLlama-13b-Instruct-hf [^4] https://huggingface.co/meta-llama/CodeLlama-13b-Instruct-hf.
@@
Dataset. As mentioned in Section 3, four-fifth of the instances from SoAyBench are utilized here to train the SoAyLLaMA. We simply take the academic query as an input and the relevant solution.
@@
Training Details. We train 13B models on 8 * NVIDIA A100 GPU (80GB) for 13 hours to accomplish a convergent loss, with a batch size of 8. 7B models are trained on 4 * NVIDIA A100 GPU (80GB) for 7 hour with a batch size of 16. All the training details are saved in the training log in our Github Repo [^5] https://anonymous.4open.science/r/SoAy-C025/readme.md.
Appendix C SoAyBench
@@
To assess AMiner API-using capabilities, the foundational AMiner APIs for LLMs to invoke and a test set should be provided. However, given the dynamic and specialization nature of academic data, with academic entities-related information rapidly changing, maintaining a test set with static answers proves challenging. To address this challenge, we clone AMiner’s APIs at a specific point in time to create a static service, from which we create a corresponding static test set including 792 meaningful academic questions (one-fifth of the whole SoAyBench) and ground truth conducted from the cloned APIs. We extract a subset of the comprehensive AMiner database, which includes over 66,000 entities such as persons and publications, at a specific point in time, September 23, 2023. This subset serves as our experimental evaluation environment, and we clone an API Service based on it. Consequently, the results returned by these APIs, given the same parameters, will remain consistent over time. This dataset and the API service provide a foundation for our subsequent construction of the test set.
Part II. Experiments Details
In this section, we will present more details related to the experiments. To verify the efficiency and effectiveness of SoAy in real applications, it is insufficient to rely solely on its performance on SoAyBench, so we designed an online evaluation to test SoAy’s performance in real-use scenarios. Beyond the research questions in Section 7, more interesting insightful conclusions can be found in the results on SoAyBench. Furthermore, to validate the universality of the SoAy method, rather than it only working for the AMiner system, we implemented SoAyGPT in two other academic information seeking scenarios and demonstrated its results in a case study manner.
Appendix D Online Evaluation
Figure 4: Distribution of questions used for online human evaluation. Questions are categorized by 1, 2, and 3-hop types, further divided into exact and ambiguous categories. The ambiguous category is specific to this online evaluation setting.
We examine the Spearman correlation coefficients among the voting results of the 15 human annotators, as shown in Figure 5. The released SoAyBench contains the templated questions derived from the AMiner SoAPI, thus testing a method’s ability to understand and utilize the SoAPIs to a certain extent. However, in real-world deployment scenarios, the goal of a good method should be to fulfill user needs more comprehensively and offer a superior user experience. To achieve this, we implement the proposed SoAy as an online application, gather 52 real user demands from the logs, and invite 10 annotators to conduct human evaluation.
Figure 5: Spearman correlation among 15 annotators
Settings. We deploy SoAy with GPT-3.5-16k as its backbone online, due to the superior speed of the 3.5 version, enhancing user experience. Over time, we gather 52 authentic questions from the application logs. The distribution of these questions is illustrated in Figure 4, categorized not only by their 1, 2, and 3-hop nature but also classified as either exact or ambiguous. Ambiguous questions, such as requests for “comparisons between two scholars”, lack standardized answers and thus yield diverse responses. We find that the majority of these 52 questions could be resolved within two hops.
To evaluate the answers, we invite five academic experts to provide responses based on their expertise, supplemented by the AMiner system and general search engines. The online system’s answers, responses directly from ChatGPT, and the aforementioned human-provided answers constitute the comparison methods.
For a balanced and fair assessment of the experimental results, we develop an annotation platform and create a detailed annotation manual. This platform allows annotators to shuffle and anonymize the answers from the three methods before each voting session. A team of 10 annotators is recruited to assess the answers provided by each method for every question across a two-day period. Their focus is on evaluating the accuracy and readability of the responses, culminating in a vote for the most effective response among the three. We also open-source the platform and its associated manual.
Evaluation Results.
Figure 6: Voting results for authentic online questions.
After identifying the three annotators with the lowest correlation to others, we exclude their input. For the remaining annotators, a majority voting strategy is employed to consolidate the votes for the three methods. The summarized results are presented in Figure 6. These results show a preference for the SoAy method’s responses in real-world scenarios across various question types, outperforming both direct ChatGPT and human expert responses.
Notably, SoAy’s responses are particularly favored for questions with exact answers. However, SoAy’s advantage is less pronounced for ambiguous questions when compared to baseline responses. This observation aligns with SoAy’s design and the assessment objectives of SoAyBench, both of which prioritize precise answering through the integration of SoAPI Library resources. Nevertheless, there’s a noticeable gap between the nature of real-world inquiries and the templated questions in SoAyBench, posing a challenge for SoAy in addressing certain scenarios. This finding points towards the potential for integrating SoAPI calls with model-intrinsic functions for a more robust response system.
Appendix E Insights on SoAyBench
E.1 Intriguing phenomena
Table 3 reveals several other intriguing phenomena: First, while SoAy excels with GPT-4, its advantages are less marked on GPT-3.5-turbo and GPT-3.5-16k (hereafter collectively referred to as GPT-3.5 backbones), trailing behind the GPT-DFSDT method with the same GPT-3.5 backbones. Notably, the DS and WS metrics for SoAy with GPT-3.5 backbones are quite high. This can be attributed to GPT-4’s superior instruction-following proficiency, which amplifies the impact of SoAy’s complex prompts on the model adept at instruction following. Second, SoAy methods exhibit superior performance on the two-hop questions than one and three-hop questions, particularly with GPT-3.5 backbones. This is likely because the solution parsing prompts sampled from the SoAyBench are biased towards two-hop questions (refer to Table 2 for more information), causing the GPT-3.5 backbones to default to two-hop solutions within the SoAy framework. On the contrary, GPT-4’s robust instruction-following ability enables it to perform consistently across all question types. Third, GPT-DFSDT exhibits a stronger understanding of one-hop questions when using GPT-3.5 backbones. It breaks down complex questions into simpler one-hop questions, determining the appropriate APIs and their parameters step by step, which enhances its comprehension of one-hop questions.
E.2 Variances
Notably, the data fluctuations marked in Table 3 do not represent variances represented by repeating the identical experiment on identical samples. Given that the $768$ test samples in SoAyBench originate from $44$ template questions, and each template is filled with different entities to create unique question instances, we compute each metric within a group of instantiated questions from the $44$ question templates. We then calculate the mean and variance across groups of different instantiated questions. Our observations indicate that SoAyGPT-4 and SoAyLLaMA delivers a more consistently stable performance across various error types, which benefits from GPT-4’s enhanced generalization capacity across a variety of entities and sufficient and diverse training data from SoAyBench.
E.3 Detailed Error Analysis
Notably, We do not rectify network errors because SoAy aims to apply the system in real-world use, implying that it is essential to test each model’s stability. During experiments, we attempt to request the models up to 3 times. If all 3 attempts fail, the situation is classified as an EE.
E.3.1 ToolLLaMA (7B)
Shows the highest Execution Error (EE) rates across all question types, which might indicate issues with the model’s ability to generate executable programs consistently. It struggles particularly with one-hop questions, which have the lowest Exact Match (EM) rate and high Wrong Solution (WS) errors. Two-hop and three-hop questions reflect similar challenges with WS and EE, although DS errors decrease slightly as the complexity increases. High variability in error rates suggests that the model might benefit from training focused on error reduction techniques. The overall score of 16.72 is the lowest among the models, reflecting significant room for improvement.
E.3.2 GPT-DFSDT
Version 3.5: One-hop questions see extremely high DS rates but lower WS and EE, suggesting issues with solution path diversity but not necessarily with execution. Two-hop and three-hop configurations show a balance of errors across DS, WS, and EE, with a slightly better EM rate for three-hop.
Version 3.5-16K: This model shows improved performance over the standard 3.5 version, with particularly notable improvements in EM and ACC for one-hop questions. DS errors are significantly reduced compared to the 3.5 version, indicating enhanced solution alignment.
Version 4: The most advanced version, showing the best performance with notably high scores, especially in one-hop questions which almost reach perfect EM and ACC. Lower WS rates across all configurations, indicating robustness in generating correct solutions.
E.3.3 SoAyGPT
Version 3.5: Generally strong in EM and ACC, particularly for one-hop questions, suggesting effective solution generation and execution. Lower EE than ToolLLaMA, indicating more reliable program execution.
Version 3.5-16K: Consistently high EM rates across question types, with particularly low DS errors in two-hop questions. Shows good balance across all error types, with no extreme weaknesses.
Version 4: Outstanding performance in one-hop questions with nearly no errors and very high EM and ACC. Two-hop and three-hop questions also show very high scores, making this version highly effective and reliable.
E.3.4 SoAyLLaMA
Chat-7B: Exhibits nearly perfect performance in one-hop questions with zero DS, WS, and WP errors. Three-hop questions maintain high performance but see a slight increase in WP errors, which still remain low compared to other models.
Code-7B: Solid performance with minimal error rates across all question types. A slight uptick in WP errors in two-hop questions could be a focal point for model refinement.
Code-13B: The best-performing model across all metrics, with impressive EM and ACC in one-hop questions and consistent strength in multi-hop configurations. Nearly negligible error rates, setting a high standard for large language models.
Appendix F Other Scenarios
F.1 SoAy implemented on CrossRef
F.1.1 Introduction
Crossref is a registration agency for DOIs (Digital Object Identifiers) established by a consortium of renowned global publishers. It registers DOI numbers for newly published and previously published journal articles. All works registered with Crossref by various publishers automatically establish citation links with each other, facilitating readers to navigate journals and access the websites of different publishers, thus significantly increasing the chances of journal articles being accessed and cited. Crossref’s open reference linking system establishes citation links to journal articles between different publisher websites. For developers, Crossref provides rich open-source API interfaces on GitHub [^6] https://github.com/CrossRef/rest-api-doc?tab=readme-ov-file, such as search and other functionalities.
F.1.2 Implementation
The codes, prompt, and APIs of this part of the implementation are available at our GitHub repo: https://github.com/RUCKBReasoning/SoAy/tree/master/scenarios/CrossRef. As shown above, we need to finish two primary stages: the offline solution library preparation stage and the online solution-based inference stage.
Figure 7: Illustration of Crossref SoAPI Library, SoAPI Graph and solution library.
API Library. As shown in Figure 7, we use 6 APIs of Crossref. The phenomenon of API coupling is prevalent among these SoAPIs. For example, the parameter doi for getWorksByDoi(·) must first be obtained by searchWorkByTItle(·). This necessitates a specific execution order where searchWorkByTItle(·) must be used before getWorksByDoi(·), and a specific execution composition where if getWorksByDoi(·) is used, searchWorkByTItle(·) is definitely used.
Solution Library Construction. The solution library is the cornerstone of the entire framework. Just as what we did for AMiner, we finish the solution library for Crossref, proving SoAy’s usability. The initial stage of SoAy involves analyzing Crossref’s SoAPI and constructing the SoAPI Graph. Following this, a process of sampling and filtering paths is applied to the SoAPI Graph, resulting in the solution library. Based on it, the prompt library is then constructed.
To better understand the performance of Soay on Crossref, let’s take a detailed look at Crossref’s prompts and cases on Figure 10
General Prompt Drafting. The General Prompt outlines a structured approach to crafting programs that utilize a set of APIs to address various queries related to literature and authors. It starts by presenting the APIs that will be employed, such as ‘searchBook‘, ‘getBook‘, and others, which are instrumental in searching for books, obtaining book details, and accessing author information. The prompt then proceeds to illustrate the process through a series of examples. The first example guides the user on how to use the ‘searchBook‘ API to find publishers in a particular subject area (XXX), iterating through the results to compile a list of publisher names. This demonstrates the initial step of narrowing down the scope of the query to a more manageable list. Following that, the second example builds upon the first by not only identifying publishers in a specific field but also by delving deeper to list articles or works associated with those publishers. This is achieved by chaining the ‘searchPublisherBySubject‘ API to ‘getPublisherBasicInfo‘ and then to ‘getPublisherWorks‘, showcasing a more complex use of the APIs to extract detailed information. The prompt concludes with a user query placeholder, encouraging the LLMs to apply the learned methods to formulate a solution. This involves selecting the correct sequence of APIs to retrieve the necessary information in response to the query. The prompt is designed to be a comprehensive guide that adapts to different levels of complexity in the user’s requests, from simple to more advanced searches.
F.1.3 Case Study
Case 1: Mathematics Publishers. The first case presents a query for a list of publishers in the mathematics field. The solution uses the ‘searchPublisherBySubject‘ API with ’math’ as the input to extract a list of publishers, along with the number of DOIs they’ve published. The result provides a collection of publisher names and their respective DOI counts, highlighting their impact in the field of math.
Case 2: DOI Numbers in Computer Science. The second case addresses the request for current DOI numbers from publishers in computer science. By utilizing ‘searchPublisherBySubject‘ to identify relevant publishers, followed by ‘getPublisherBasicInfo‘ to obtain DOI counts, the solution compiles a list of publishers with their current DOI numbers, showcasing their recent publication activities.
Case 3: Article Titles in Computer Science. The third case involves listing articles in the computer science domain. It starts with ‘searchPublisherBySubject‘ to find publishers, then uses ‘getPublisherBasicInfo‘ and ‘getPublisherWorks‘ to retrieve a list of article titles. The output is a concise list of article titles that represent the range of research within computer science.
F.2 SoAy implemented on OpenLibrary
F.2.1 Introduction
Open Library is an online project aimed at creating a comprehensive database of all published books. Operated by the Internet Archive, the website has been serving the public since 2006. It provides information pages for thousands of books, displaying their summaries, covers, publication details, and more. Additionally, it features personal pages for numerous authors, detailing their biographical information, major works, and areas of expertise. For developers, the website also offers web APIs such as search functionality, which are documented in the Public APIs project.
F.2.2 Implemention
The codes, prompt, and APIs of this part of the implementation are also available at our GitHub repo: https://github.com/RUCKBReasoning/SoAy/tree/master/scenarios/OpenLibrary. The same as AMiner and CrossRef, we need to finish two primary stages: the offline solution library preparation stage and the online solution-based inference stage.
Figure 8: Illustration of OpenLibrary SoAPI Library, SoAPI Graph and solution library.
API Library. As shown in Figure 8, we use 6 APIs of Crossref. The phenomenon of API coupling is prevalent among these SoAPIs. For example, the parameter subject for searchSubject(·) must first be obtained by getBook(·) or getAuthorBasicInfo(·). This necessitates a specific execution order where getBook(·) or getAuthorBasicInfo(·) must be used before searchSubject(·), and a specific execution composition where if searchSubject(·) is used, getBook(·) or getAuthorBasicInfo(·) is definitely used.
Solution Library. The solution library is the cornerstone of the entire framework. Just as what we did for AMiner, we finish the solution library for OpenLibrary, proving SoAy’s usability. The initial stage of SoAy involves analyzing OpenLibrary’s SoAPI and constructing the SoAPI Graph. Following this, a process of sampling and filtering paths is applied to the SoAPI Graph, resulting in the solution library. Based on it, the prompt library is then constructed.
General Prompt. To better understand the performance of Soay on OpenLibrary, let’s take a detailed look at OpenLibrary’s prompts and cases on Figure 11 It starts by presenting the APIs that will be employed, such as ‘searchBook‘, ‘getBook‘, and others, which are instrumental in searching for books, obtaining book details, and accessing author information. The prompt then proceeds to illustrate the process through a series of examples. The first example guides the user on how to use the ‘searchBook‘ API to find publishers in a particular subject area (XXX), iterating through the results to compile a list of publisher names. This demonstrates the initial step of narrowing down the scope of the query to a more manageable list. The same principle applies to the two other examples. The prompt concludes with a user query placeholder, encouraging the LLMs to apply the learned methods to formulate a solution. This involves selecting the correct sequence of APIs to retrieve the necessary information in response to the query. The prompt is designed to be a comprehensive guide that adapts to different levels of complexity in the user’s requests, from simple to more advanced searches.
F.2.3 Case Study
Case 1: Identifying Mathematics Publishers. In the first case, the objective is to identify publishers that specialize in mathematics. The process begins by utilizing the ‘searchPublisherBySubject‘ API, specifying ’math’ as the subject parameter. This call returns a list of publishers active in the field of mathematics, each accompanied by the number of DOIs (Digital Object Identifiers) they have generated. The output is a curated list of publisher names alongside their DOI counts, which serves as an indicator of their contributions and presence in the mathematical literature space.
Case 2: Obtaining DOI Numbers for Computer Science Publishers. The second case focuses on obtaining the current DOI numbers for publishers within the computer science field. The ‘searchPublisherBySubject‘ API is employed to first filter out publishers related to computer science. Once a list of these publishers is obtained, the ‘getPublisherBasicInfo‘ API is then used to fetch the up-to-date DOI numbers for each publisher. The result is a compilation of publisher names paired with their current DOI counts, providing a snapshot of their recent scholarly output and engagement within the field.
Case 3: Listing Computer Science Articles. The third case aims to extract a list of article titles from the computer science domain. The process initiates with the ‘searchPublisherBySubject‘ API to pinpoint publishers in the field, followed by the use of ‘getPublisherBasicInfo‘ to acquire the publisher’s unique ID. With the publisher ID in hand, the ‘getPublisherWorks‘ API is then applied to retrieve a comprehensive list of article titles associated with the publisher. The final output is a succinct array of article titles, offering a glimpse into the diverse range of research topics and findings published within the realm of computer science.
Part III. Prompts
In this section, we will comprehensively present the prompts used in the data generation process of SoAy, as well as the implementation process of SoAyGPT on AMiner and two other scenarios.
Figure 9: A Preview of the prompt library. $P_{s}$ signifies the prompt for solution retrieval, $P_{p}$ represents the prompt for program generation, and $P_{a}$ denotes the prompt for answer generation. All the black text except the ellipsis is the original prompt content, the yellow highlighted text is the annotations added for better understanding in the context of creating a prompt, and the purple highlighted text is the content waiting for the LLM output.
Figure 10: A Preview of the prompt library and three cases for Crossref.
Figure 11: A Preview of the prompt library and three cases for OpenLibrary.
Figure 12: A Preview of the prompts used in the data preparation process.
Figure 13: The screenshots of the demonstration systems https://soay.aminer.cn and online service widget https://chatglm.cn/main/gdetail/65bf5a99396389a73ace9352 powered by SoAy.

### v2

@@ 8. Conclusion
In this work, we propose an API-using methodology, SoAy, to apply LLMs to academic information searching. SoAy generates a solution firstly, which is a sequence of APIs, and then uses this solution to guide the code generation process. By parsing the solution, SoAy can handle the coupling relation between a set of APIs. Code generation and execution process makes SoAy more efficient than previous methods using multi-step reasoning. The core technique of SoAy is the high-quality automatic built dataset, SoAyBench, providing 3,960 academic triplets data. Based on the training set of SoAyBench, we build SoAyLLaMA and SoAyGPT, separately using the tuning method and tuning-free method. To evaluate the AMiner API-using capability of different models and methods, we propose a comprehensive evaluation strategy, SoAyEval. Experiments show SoAy’s effectiveness, efficiency, and usability. SoAy has been applied to real-world service and has been accessed more than 54,800 times. The implementation of SoAy to various applications in academic domain highlights the usability of SoAy. In the future, we will continue to enlarge the SoAyBench and apply SoAy to more scenarios and domains. In the future, we will continue to enlarge the SoAyBench and enrich the queries in SoAyBench based on the real-world application user logs.
Acknowledgements.
@@
This work is supported by the National Key Research & Develop Plan (2023YFF0725100) and National Natural Science Foundation of China (62322214, U23A20299, U24B20144, 62172424, 62276270). This project would like to thank the data annotators at Zhipu AI for their help and support. All GPU compute and API costs are provided by Zhipu AI.
@@
Appendix A Application of SoAy
As of the submission date, SoAy has been deployed for 10 months and has been accessed more than 54,800 times in the real-world application at https://www.aminer.cn/ and https://chatglm.cn/main/gdetail/65bf5a99396389a73ace9352, providing users with accurate academic information and helping them stay informed with up-to-date insights throughout their research process.
Appendix B SoAyGPT
@@
SoAyGPT is an alignment method designed for a series of large-scale models that are closed-source yet possess high instruction following capabilities. We implement SoAyGPT on three distinct versions of GPT, in particular 3.5: gpt-3.5-turbo-0613, 3.5-16k: gpt-3.5-turbo-16k-0613, and 4: gpt-4-0613. Specifically, SoAyGPT employs in-context learning to construct three LLM agents (Wang et al., 2024) (Solution Generation Agent, Code Generation Agent, and Answer Generation Agent) and an environment operable by these agents to complete the alignment process.
Identifying the solution first and then generating the code is less challenging for LLMs, compared to directly code generation, reducing the likelihood of producing incorrect answers. Additionally, the advantage of generating the solution and code separately, rather than simultaneously, is that each step allows for more information and examples to be provided within a limited input space.
The agents and environments are designed as follows. Prompts for each agent are shown in Figure 7.
Appendix C SoAyLLaMA
@@
SoAyLLaMA is designed for supervised fine-tuning of open-source models when training resources are available. During the training stage, SoAyLLaMA uses the API description and the query as input for each data sample, and combines the solution and code as the response. During the inference stage, SoAyLLaMA concurrently completes the Planning and Formatting processes to derive solutions and code. Similar to SoAyGPT, it utilizes the environment to execute and obtain answers based on $K$. We conduct the fine-tuning on 3 different variants of LLaMA Model, respectively Chat-7B: Llama-2-7b-chat-hf [^1] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf, Code-7B: CodeLlama-7b-Instruct-hf [^2] https://www.modelscope.cn/models/AI-ModelScope/CodeLlama-7b-hf/summary, Code-13B: CodeLlama-13b-Instruct-hf [^3] https://huggingface.co/meta-llama/CodeLlama-13b-Instruct-hf.
@@
Dataset. As mentioned in Section 3, four-fifth of the instances from SoAyBench are utilized here to train the SoAyLLaMA.
@@
Training Details. We train 13B models on 8 NVIDIA A100 GPUs (80GB) for 13 hours to reach a convergent loss, using a batch size of 8. Additionally, we train 7B models on 4 NVIDIA A100 GPUs (80GB) for 7 hours with a batch size of 16. All the training details are saved in the training log in our Github Repo [^4] https://anonymous.4open.science/r/SoAy-C025/readme.md.
Appendix D SoAyBench
@@
To assess AMiner API-using capabilities, the foundational AMiner APIs for LLMs to invoke and a test set should be provided. However, given the dynamic and specialization nature of academic data, with academic entities-related information rapidly changing, maintaining a test set with static answers proves challenging. To address this challenge, we clone AMiner’s APIs at a specific point in time to create a static service, from which we create a corresponding static test set including 792 meaningful academic questions (one-fifth of the whole SoAyBench) and the corresponding ground truth conducted from the cloned APIs. We extract a subset of the comprehensive AMiner database, which includes over 66,000 entities such as persons and publications, at a specific point in time, September 23, 2023. This subset serves as our experimental evaluation environment, and we clone an API Service based on it. Consequently, the results returned by these APIs, given the same parameters, will remain consistent over time. This dataset and the API service provide a foundation for our subsequent construction of the test set.
Appendix E Implementation on Other Scenarios
Although SoAy is developed for the AMiner API, it can be seen as a general method for enabling large language models (LLMs) to use domain-specific APIs. To verify this generality, and also to prove that similar coupling phenomena, like those observed with the AMiner API, also exist in other domain-specific APIs, we deploy SoAy in two other scenarios that provide information-seeking APIs, OpenLibrary and CrossRef. Due to GPU resource limitations, we did not train separate models for these two new scenarios. Instead, we implemented corresponding SoAyGPT models to validate the feasibility of setting up SoAy in a new API domain. Details can be found at https://github.com/RUCKBReasoning/SoAy/tree/master/scenarios.
Figure 7. The prompt is designed during the inference of SoAyGPT. Specifically: $P_{s}$ signifies the prompt for solution generation, which instructs the LLMs to generate a solution from the user’s natural language query. $P_{p}$ represents the prompt for code generation. This prompt is used to instruct the LLMs to generate code from the user’s query and the corresponding solution. Note that the code generation prompt is tailored for each solution in the solution library $L_{S}$. If the given solution is found in $L_{S}$, its specific $P_{p}$ is used; otherwise, a general $P_{p}$ is employed. This general prompt is composed by sampling (Query, Solution, Code) triplets from the training data for each solution in $L_{S}$. $P_{a}$ denotes the prompt for answer generation, used to instruct the LLM to generate a natural language answer based on the query and the code execution result. In the prompt content: All black text, except the ellipses, represents the original prompt content. Orange highlighted text includes annotations added for better understanding in the context of creating a prompt. Purple highlighted text indicates the content awaiting LLM output.
[…]
The prompt is designed during the inference of SoAyGPT.